OpenStudy (anonymous):

ABCD is a parallelogram. If AB =2x - y, CD = 7, BC = x + y , AD = 2 find x and y

OpenStudy (abmon98):

Opposite sides of a parallelogram are the same length (congruent).

OpenStudy (abmon98):

2x-y=2 x+y=7 y cancel out each other if the signs are opposite in a simultaneous equation we add 3x=9 x=3 subsititute x=3 2(3)-y=2 -y=-4 y=4